OpenStudy (anonymous):

How do I solve for c & H using elimination so far I have C=200+50H/ C=300+25H help me please

OpenStudy (anonymous):

C=200+50H C=300+25H 200+50 = 300+25H can you go from there?

OpenStudy (anonymous):

typo: 200+50H = 300+25H

OpenStudy (anonymous):

Use a system of equations. (Set both equations equal to each other, since the both equal C) (like MathMind has shown). Then you take the answer for H and plug it back into one of the equations for C!

OpenStudy (anonymous):

do I add 200 w/ 300 and 50 w/ 25 then divide

OpenStudy (anonymous):

hmm not quite: 200-300 = 25H-50H -100 = -25H

OpenStudy (anonymous):

so Isubtract both sides

OpenStudy (anonymous):

so is H=0.25

OpenStudy (anonymous):

yup. Even thou adding them comes up to the same answer i'm not sure if it'll always work 200+300=25H+50H 500=75H (simplify by dividing everything by 5) 100 = 25H

OpenStudy (anonymous):

indeed. H=0.25, now just plug it back into any of those C eqt. and yu can find C

OpenStudy (anonymous):

so take 100 divded by 25 to get C

OpenStudy (anonymous):

I got 6.6666666

OpenStudy (anonymous):

hmm.. not quite.. C=200+50H C=300+25H with H = 100/25 = 4

OpenStudy (anonymous):

for H I got H=0.25

OpenStudy (anonymous):

Now I am tryn to solve for C so don't I divide again

OpenStudy (anonymous):

your H is wrong. 100 = 25H (devide both sides by 25) 100/25 = H 4 = H You have the equation for C to be: C=200+50H C=300+25H if you know H = 4: C=200+50(4) = 400 C=300+25(4) = 400

OpenStudy (anonymous):

so I start by subtracting then divide to solve for h then to solve for C I multiply both sides by 4 which made them equal.

OpenStudy (anonymous):

for C all you have to do is substitute whatever value you found for H in the C equation

OpenStudy (anonymous):

ok I see so tht is how they become even

OpenStudy (anonymous):

I think I got it thank you when I get to the other ones I will be back lol

OpenStudy (anonymous):

ok how do you check for accuracy?

OpenStudy (anonymous):

C=200+50(4) = 400 C=300+25(4) = 400 that's your check

OpenStudy (anonymous):

O so c=4 h=400 and that is how I prove it

OpenStudy (anonymous):

nope.. what value did you find for H after you solve everything?

OpenStudy (anonymous):

4 right

OpenStudy (anonymous):

or do I add 400 wth 400

OpenStudy (anonymous):

right, H = 4 we also have two equations: C=200+50(H) C=300+25(H) If C are equal to those two equations than it's right to say: 200+50(H)=300+25(H) Both sides should be equal after we plug in H=4 200+50(4)=300+25(4) 400 = 400 . Yay! It's true!

OpenStudy (anonymous):

ok so umm I think I have it maybe

OpenStudy (anonymous):

hmm ok. If yu still have a question just ask

OpenStudy (anonymous):

ok I will I am on another problem now

OpenStudy (anonymous):

asking me a loan amout? and loan amount at the end of 5 years? and the unpaid balance at the end of 10 years I think I remember this I am going to try then show u

OpenStudy (anonymous):

ok I am lost and confused I thought it was PMT I have 250.000 6% interest and 30 years it asked me what are th emonthly payments